# Distributed Model Checking in Graphs Classes

### v1

@@ 3.2 Graphs of Bounded Treedepth
We stress the fact that a decomposition forest $F$ of $G$ must be a subgraph of $G$, which is very convenient for distributed algorithms because this enables us to communicate along the edges of $F$. There exist graphs of treedepth $t$, in which any decomposition forest is of depth $\Omega(2^{t})$, for example the paths on $12^{t}-$ 1 vertices. Nevertheless, graphs of treedepth at most $t$ have paths of at most $12^{t}-$ 1 vertices. In particular, for any class of graphs of bounded treedepth, all possible decomposition forests are also of constant (although exponentially larger) depth.
Proposition 3 ([67]).
Any decomposition forest of a graph of tree-depth at most $t$ belongs to $\mathcal{F}_{2^{t}}$.
In the following lemma, we show that any distributed algorithm aiming at cheching a quantifier-free formula in a treedepth-$t$ induced subgraph $H$ of a labeled graph $G$, may rather check an equivalent formula on a decomposition forest of $H$.
Lemma 6.
Let $\Lambda$ be a finite set of labels, and let $t>0$. For every quantifier free formula $\varphi(\overline{x})\in\text{FO}[\Lambda]$, there exists a finite set $\widehat{\Lambda}$ depending only on $t$ and $\Lambda$, and an lca-reduced formula $\widehat{\varphi}(\overline{x})\in\text{FO}[\widehat{\Lambda}]$ depending in $\varphi$, $\Lambda$ and $t$, such that, for every $\Lambda$-labeled treedepth-$t$ induced subgraph $(H,L)$ of a $\Lambda$-labeled graph $(G,L)$, and for every decomposition forest $F$ of $H$, there exists a $\widehat{\Lambda}$-labeling $\widehat{L}$ of $G$ such that
$\textsf{true}(\varphi,H,L)=\textsf{true}(\widehat{\varphi},F,\widehat{L}).$
Moreover, there is a 1-round CONGEST algorithm that computes $\widehat{\Lambda}$, $\widehat{\varphi}$, and $\widehat{L}$. The algorithm assumes that each node $u$ initially knows $\Lambda$, $\varphi$, $L(u)$, $t$, whether it belongs to $H$, as well as its depth and its incident edges in $F$, and returns $\widehat{\Lambda}$, $\widehat{\varphi}$, and $\widehat{L}(u)$.
@@
Let $d=2^{t}$, and let us consider a $\Lambda$-labeled graph $(G,L)$, a treedepth-$t$ induced subgraph $H$ of $G$, and a decomposition forest $F$ of $H$. The idea of the construction is the following. Each node $u$ of $F$ can be labeled by a subset of $[d]$ indicating the set of levels of $F$ containing a neighbor of $u$ in $H$. Then, the $\textsf{adj}(x,y)$ predicate can be transformed into label and lca predicates on $x$ and $y$. More precisely, for every node $v\in V(H)$, let $\textsf{Levels}(v)\subseteq[d]$ as the set of levels of $F$ containing ancestors $u$ of $v$ such that $\{u,v\}\in E(H)$. Then, let $\widehat{\Lambda}=\Lambda\cup[d]$ (we assume w.l.o.g. that $\Lambda\cap[d]=\emptyset$), and let us consider the $\widehat{\Lambda}$-labeling $\widehat{L}$ of $G$ defined by
$\widehat{L}(v)=\begin{cases}L(v)\cup\textsf{Levels}(v)&\text{if }v\in V(F),\\ L(v)&\text{otherwise.}\end{cases}$
@@
For each $i\in[d]$, we denote $\textsf{level}_{i}(x)$ the predicate $\textsf{lab}_{i}(x)$, which is true when $x$ has a neighbor at depth $i$. Finally, we define $\widehat{\varphi}$ by replacing each $\textsf{adj}(x,y)$ predicates by
$\bigvee_{1\leq i<j\leq d}\Bigg{[}\bigg{(}\textsf{lca}_{i}(x,x)\wedge\textsf{lca}_{j}(y,y)\wedge\textsf{level}_{j}(x)\bigg{)}\vee\bigg{(}\textsf{lca}_{j}(x,x)\wedge\textsf{lca}_{i}(y,y)\wedge\textsf{level}_{j}(y)\bigg{)}\Bigg{]}.$
@@
and every equality predicate by lca predicates as we explained in Lemma 1. This way we obtain a lca-reduced formula $\widehat{\varphi}$ such that $\textsf{true}(\varphi,H,L)=\textsf{true}(\widehat{\varphi},F,L)$.
@@
Finally, observe that every node $u$ can compute $\widehat{\Lambda},\widehat{\varphi}$, and $\widehat{L}(u)$ in a single communication round. In fact $\widehat{\Lambda},\widehat{\varphi}$, and $\widehat{L}(u)$ for every $u\in V(G)\setminus V(H)$ are computed without any communication, as these values do not depend on the instance. To compute $\widehat{L}(u)$ for $u\in V(F)$, each node broadcasts its depth in $F$ to all its neighbors. After this communication round, $\textsf{Levels}(u)$ can be set as the set of neighbors with a smaller depth in $F$, and the computation of $\widehat{L}(u)=L(u)\cup\textsf{Levels}(u)$ follows. ∎
@@ 3.3 Graphs of Bounded Expansion
Let us fix a graph class $\mathcal{G}$ of bounded expansion $f$. For a set of labels $\Lambda$, let $\mathcal{G}[\Lambda]$ be the set of $\Lambda$-labeled graphs $(G,L)$ such that $G\in\mathcal{G}$. Let $(G,L)\in\mathcal{G}[\Lambda]$ be the input instance. For every integer $p$, let $C_{p}$ be a $(p,f(p))$-treedepth coloring of $G$, and let $\textsf{Col}(p)$ be the collection of sets $U\subseteq[f(p)]$ with cardinality at most $p$.
3.3.1 Skeletons and Compatible Labelings
For every set $U\in\textsf{Col}(p)$, let $F^{U}$ be a decomposition forest of the subgraph $G[U]$ of $G$ induced by the set of nodes colored with colors in $U$. By definition of $C_{p}$, the graph $G[U]$ is of treedepth at most $p$. Hence, by Proposition 3, $F^{U}\in\mathcal{F}_{2^{p}}$. The set
$\tau_{p}=\{F^{U}\mid U\in\textsf{Col}(p)\}$

### v2

@@ 3.2 Quantifier Elimination on Graphs of Bounded Treedepth
We stress the fact that a decomposition forest $F$ of $G$ must be a subgraph of $G$, which is very convenient for distributed algorithms because this enables communication along the edges of $F$. Note that there exist graphs of treedepth $t$ in which any decomposition forest is of diameter $\Omega(2^{t})$. This is for example the case of paths on $12^{t}-$ 1 vertices. Nevertheless, graphs of treedepth at most $t$ have paths of at most $12^{t}-$ 1 vertices. In particular, for any class of graphs of bounded treedepth, all possible decomposition forests also have trees of constant (although exponentially larger) diameter.
Proposition 3.10 ([54]).
Any decomposition forest of a graph of treedepth at most $t$ belongs to $\mathcal{F}_{2^{t}}$.
In the following lemma, we show that any distributed algorithm aiming at checking a quantifier-free formula in a treedepth-$t$ induced subgraph $H$ of a labeled graph $G$, may rather check an equivalent formula on a decomposition forest of $H$.
Lemma 3.11.
Let $\Lambda$ be a finite set of labels, and let $t>0$. For every quantifier free formula $\varphi(\overline{x})\in\mathrm{FO}[\Lambda]$, there exists a finite set $\widehat{\Lambda}$ depending only on $t$ and $\Lambda$, and an lca-reduced formula $\widehat{\varphi}(\overline{x})\in\mathrm{FO}[\widehat{\Lambda}]$ depending on $\varphi$, $\Lambda$, and $t$, such that, for every $\Lambda$-labeled treedepth-$t$ induced subgraph $(H,\ell)$ of a $\Lambda$-labeled graph $(G,\ell)$, and for every decomposition forest $F$ of $H$, there exists a $\widehat{\Lambda}$-labeling $\widehat{\ell}$ of $G$ such that
$\textsf{true}(\varphi,H,\ell)=\textsf{true}(\widehat{\varphi},F,\widehat{\ell}).$
Moreover, there is a 1-round CONGEST algorithm that computes $\widehat{\ell}$, assuming each node $v$ is given as input $\ell(v)$, whether it belongs to $F$ and, if applicable, its depth and parent in $F$.
@@
Each node $u$ initially knows $\Lambda$, $\varphi$, $\ell(u)$, $t$, whether it belongs to $H$, and, if applicable, its depth and parent in $F$. Node $u$ returns $\widehat{\Lambda}$, $\widehat{\varphi}$, and ^$\widehat{\ell}(u)$. Let $d=2^{t}$, and let us consider a $\Lambda$-labeled graph $(G,\ell)$, a treedepth-$t$ induced subgraph $H$ of $G$, and a decomposition forest $F$ of $H$. The idea of the construction is the following. Each node $u$ of $F$ can be labeled by a subset of $[0,d-1]$ indicating the set of levels of $F$ containing a neighbor of $u$ in $H$. Then, the $\textsf{adj}(x,y)$ predicate can be transformed into label and lca predicates on $x$ and $y$. More precisely, for every node $v\in V(H)$, let $\textsf{Levels}(v)\subseteq[d]$ be the set of levels of $F$ containing ancestors $u$ of $v$ such that $\{u,v\}\in E(H)$. Then, let $\widehat{\Lambda}=\Lambda\cup[0,d-1]$ (we assume w.l.o.g. that $\Lambda\cap[0,d-1]=\emptyset$), and let us consider the $\widehat{\Lambda}$-labeling $\widehat{\ell}$ of $G$ defined by
^$\widehat{\ell}(v)=\begin{cases}\ell(v)\cup\textsf{Levels}(v)&\text{if }v\in V(F),\\ \ell(v)&\text{otherwise.}\end{cases}$
@@
For each $i\in[0,d-1]$, we denote by $\textsf{level}_{i}(x)$ the predicate $\textsf{lab}_{i}(x)$, which is true when $x$ has a neighbor at depth $i$. Finally, we define $\widehat{\varphi}$ by replacing each $\textsf{adj}(x,y)$ predicates by
$\bigvee_{1\leq i<j\leq d-1}\Bigg{[}\bigg{(}\textsf{lca}_{j}(x,x)\wedge\textsf{lca}_{i}(y,y)\wedge\textsf{level}_{i}(x)\bigg{)}\vee\bigg{(}\textsf{lca}_{i}(x,x)\wedge\textsf{lca}_{j}(y,y)\wedge\textsf{level}_{i}(y)\bigg{)}\Bigg{]},$
@@
and every equality predicate by lca predicates as we explained in Lemma 3.1. In this way, we obtain a lca-reduced formula $\widehat{\varphi}$ such that $\textsf{true}(\varphi,H,\ell)=\textsf{true}(\widehat{\varphi},F,\ell)$.
@@
Finally, observe that every node $u$ can compute $\widehat{\Lambda},\widehat{\varphi}$, and ^$\widehat{\ell}(u)$ in a single communication round. In fact $\widehat{\Lambda},\widehat{\varphi}$, and ^$\widehat{\ell}(u)$ for every $u\in V(G)\smallsetminus V(H)$ are computed without any communication, as these values do not depend on the instance. To compute ^$\widehat{\ell}(u)$ for $u\in V(F)$, each node broadcasts its depth in $F$ to all its neighbors. After this communication round, $\textsf{Levels}(u)$ can be set as the set of neighbors with a smaller depth in $F$, and the computation of ^$\widehat{\ell}(u)=\ell(u)\cup\textsf{Levels}(u)$ follows. ∎
We now prove that existential quantifier in formulas on graphs of bounded treedepth can be eliminated efficiently. The reduction results in a lca-reduced formula to be checked on a decomposition forest of the considered graph.
Lemma 3.12.
Let $\Lambda$ be a set of labels, $\varphi(\overline{x})=\exists\overline{y},\zeta(\overline{x},\overline{y})\wedge\mathrm{ball}((\overline{x},\overline{y}),x_{1},r)$ be a $r$-local existential formula in $\mathrm{FO}[\Lambda]$, where $r\in\mathbb{N}$ and $\zeta$ is a quantifier free formula. There exists a set of labels $\widehat{\Lambda}$, and an lca-reduced formula $\widehat{\varphi}(\overline{x})$ such that, for every treedepth-$t$ induced subgraph $(H,\ell)$ of a $\Lambda$-labeled graph $(G,\ell)$, and for every decomposition forest $F$ of $H$, there exists a $\widehat{\Lambda}$-labeling $\widehat{\ell}$ such that:
$\textsf{true}(\varphi,H,\ell)=\textsf{true}(\widehat{\varphi},F,\widehat{\ell})=\textsf{true}(\widehat{\varphi}\wedge\mathrm{ball}(\overline{x},x_{1},2^{t}),F,\widehat{\ell}).$
Moreover, there exists an algorithm computing $\widehat{\ell}$ in $\mathcal{O}(1)$ CONGEST rounds, assuming each node $v$ is given as input $\ell(v)$, whether it belongs to $F$ and, if applicable, its depth and parent in $F$.
The lemma above considers both a graph $H$ of bounded treedepth, and a decomposition forest $F$. The distances in both structures are not the same, and two vertices may be further apart in $F$ than they are in $H$. However, the connected components remain the same in both graphs.
Proof.
We are assuming that each node $u$ initially knows $\Lambda,\varphi,\ell(u)$, whether it belongs to $H$, and, if applicable, its depth and parent in $F$. We consider graphs of treedepth at most $t$, thus their connected components have diameter at most $2^{t}$. If $r<2^{t}$, then, by Lemma 2.5 we can rewrite $\mathrm{ball}(z,x_{1},r)$ as an existential formula of the form $\exists\overline{w},\psi(x_{1},z,\overline{w})\wedge\mathrm{ball}((z,\overline{w}),x_{1},2^{t})$. Thus, without loss of generality, we consider the case $r\geq 2^{t}$. Therefore, the balls of radius $r$ are then exactly the connected components of $H$.
By Lemma 3.11, there exists an lca-reduced formula $\zeta_{1}(\overline{x},\overline{y})$ such that, for every induced labeled subgraph $(H,\ell)$ of treedepth at most $t$, for every decomposition forest $F$ of $H$, there exists a labeling $\ell_{1}$ such that $\textsf{true}(\zeta,H,\ell)=\textsf{true}(\zeta_{1},F,\ell_{1})$, and $\ell_{1}$ can be computed in $\mathcal{O}(1)$ rounds. Moreover, since connected components are the same in $H$ and $F$, we also have
$\textsf{true}(\zeta\wedge\mathrm{ball}((\overline{x},\overline{y}),x_{1},r),H,\ell)=\textsf{true}(\zeta_{1}\wedge\mathrm{ball}((\overline{x},\overline{y}),x_{1},r),F,\ell_{1}).$
The proof is completed by applying Lemma 3.9 to the formula $\exists\overline{y},\zeta_{1}(\overline{x},\overline{y})\wedge\mathrm{ball}((\overline{x},\overline{y}),x_{1},r)$. ∎
@@ 3.3 Graphs of Bounded Expansion
Let us fix a graph class $\mathcal{G}$ of expansion $f$. Given a set of labels $\Lambda$, let $\mathcal{G}[\Lambda]$ be the set of $\Lambda$-labeled graphs $(G,\ell)$ such that $G\in\mathcal{G}$. Let $(G,\ell)\in\mathcal{G}[\Lambda]$ be the input instance. For every integer $p$, let $\textsf{Col}(p)$ be the collection of sets $U\subseteq[f(p)]$ with cardinality at most $p$.
3.3.1 Skeletons
We introduce a central notion for our proof of Theorem 1.1.
Definition 3.13.
Let $p\in\mathbb{N}$, and let $G$ be a graph of expansion $f$. A $p$-skeleton of $G$ is a triple $S=(G,c,\delta)$ where $c$ is $(p,f(p))$-treedepth coloring of $G$, and $\delta$ is a function mapping every $U\in\textsf{Col}(p)$ to a decomposition forest $F^{U}$ of the subgraph $G[U]$ induced by nodes with colors in $U$. Labeled $p$-skeletons are pairs $(S,\ell)$, where $S=(G,c,\delta)$ is a $p$-skeleton, and $\ell$ is a labeling of $G$.